One thousand raffle tickets are sold for $5.00 each. One grand prize of $1000 and two consolation prizes of $200 each will be awarded. Jeremy purchases one ticket. Find his expected value. Show your work for full credit.
Found 2 solutions by ewatrrr, Edwin McCravy:
Answer by ewatrrr(24785)

 
Hi
Note: in red: prize amount - the $5 cost of the ticket
One thousand raffle tickets are sold for $5.00 each.
One grand prize of $1000 and two consolation prizes of $200 each will be awarded.
total prize money awarded = $1000, $200, 200
E(X)=∑X*P(X) =
= .995 + .195 + .195 - 4.985 = -$3.60

Answer by Edwin McCravy(20086)

One thousand raffle tickets are sold for $5.00 each. One grand
prize of $1000 and two consolation prizes of $200 each will be
awarded. Jeremy purchases one ticket. Find his expected value.
Show your work for full credit.


[The other tutor forgot to subtract the price of the ticket from the
winnings, so her answer was a penny short of the expected loss of
$3.60.]

    Event                  x = winning   P(x)        xP(x)
------------------------------------------------------------
wins grand prize              $995       1/1000    995/1000
wins a consolation prize      $195       2/1000    390/1000
doesn't win                   -$5      997/1000  -4985/1000
--------------------------------------------------------------------
                                 E(x) = ΣxP(x) = -3600/1000 = -$3.60

If a raffle exactly like this one were held many times, and Jeremy bought 
one ticket every time, then he would expect to average having lost $3.60 
per time he entered.

Edwin